# Measuring AI Ability to Complete Long Tasks

### v2

@@ Abstract
Despite rapid progress on AI benchmarks, the real-world meaning of benchmark performance remains unclear. To quantify the capabilities of AI systems in terms of human capabilities, we propose a new metric: 50%-task-completion time horizon. This is the time humans typically take to complete tasks that AI models can complete with 50% success rate. We first timed humans with relevant domain expertise on a combination of RE-Bench, HCAST, and 66 novel shorter tasks. On these tasks, current frontier AI models such as Claude 3.7 Sonnet have a 50% time horizon of around 50 minutes. Furthermore, frontier AI time horizon has been doubling approximately every seven months since 2019, though the trend may have accelerated in 2024. The increase in AI models’ time horizons seems to be primarily driven by greater reliability and ability to adapt to mistakes, combined with better logical reasoning and tool use capabilities. We discuss the limitations of our results—including their degree of external validity—and the implications of increased autonomy for dangerous capabilities. If these results generalize to real-world software tasks, extrapolation of this trend predicts that within 5 years, AI systems will be capable of automating many software tasks that currently take humans a month.
@@ 1 Introduction
In the last five years, frontier AI systems have undergone a dramatic transformation in capabilities, evolving from basic text generation [1] to autonomously executing complex multi-hour machine learning research projects [2]. Sufficiently capable AIs could perform dangerous, highly complex actions like autonomous development of chemical, biological, radiological or nuclear weapons (CBRN) and self-replication and adaptation outside human control [3]. Understanding AI capabilities helps inform the development of safety guardrails as systems become increasingly powerful. In particular, many frontier AI developers have committed to using measures of specific AI capabilities to determine the necessary risk mitigations for their frontier AI systems. [^1] An updated list of frontier AI safety policies can be found at: https://metr.org/faisc. Robust benchmarks that can accurately track and forecast AI capabilities thus form the foundation for responsible AI governance and risk mitigation.
@@
However, existing benchmarks face several key limitations. First, they often consist of artificial rather than economically valuable tasks. Second, benchmarks are often adversarially selected for tasks that current models struggle with compared to humans, [^2] For example, HellaSwag [4] and Humanity’s Last Exam [5] were both generated by adversarially filtering problems against the best performing language models available at the time. biasing the comparison to human performance. Most critically, individual benchmarks saturate increasingly quickly [6], and we lack a more general, intuitive, and quantitative way to compare between different benchmarks, [^3] As with our tasks, SWE-bench Verified [7] does come with human-estimated task completion times. We use SWE-bench Verified tasks and accompanying time estimates to validate our main result in Section 6.3. which prevents meaningful comparison between models of vastly different capabilities (e.g., GPT-2 versus o1). As a consequence, while the last few years have seen dramatic increases in AI performance on many individual benchmarks, understanding the progress of AI capabilities in general has required estimating the qualitative difficulty of the latest benchmarks AI systems can pass.
@@
We propose tracking AI progress over time using the task completion time horizon: the duration of tasks that models can complete at a certain success probability, providing an intuitive measure of real-world capability compared to humans. As models may not reliably complete all tasks of a given length, we operationalize this by measuring the X%-(task completion) time horizon–the length of tasks that models can complete approximately Xof the time.
We prototype this methodology using three datasets designed to capture skills required for research or software engineering (Section 3.1), totaling 170 tasks with a wide range of difficulty: HCAST [8], RE-Bench [2], and Software Atomic Actions (SWAA), a new suite of shorter software tasks that can measure pre-2023 models (Section 3.1.3). Using skilled human baseliners, we estimate the duration that a domain knowledgeable human (without task-specific context) takes to complete these tasks (Section 3.2). We evaluate the performance of 13 frontier models from 2019 to 2025 on these tasks (Section 3.3). Using methodology inspired by human psychometric studies, we then estimate the duration of tasks that models can complete with 50% success rate—the 50% time horizon (Section 4.1).
We find that the 50% time horizon has been growing exponentially from 2019–2025 on our tasks, with a doubling time of approximately seven months (Figure 1). We compare our main result with our exploratory results on non-SWAA tasks, finding that the 2023–2025 horizon growth rate is consistent with the 2019–2025 rate to within 11% (Section 6.1). We also measure the 80% time horizon of models (Figure 6) and find a similar trend, though horizons are roughly 5x shorter. This progress appears to be driven by several key factors: improved logical reasoning capabilities, better tool use capabilities, and greater reliability and self-awareness in task execution (Section 5).  [^4] Code to reproduce our figures can be found at: https://github.com/METR/eval-analysis-public We also note several limitations of current systems—notably, performance is much lower on less structured, “messier” tasks (Section 6.2).
Since our tasks do not perfectly represent the average segment of intellectual labor by researchers and software engineers, this raises the question of external validity (Section 6): whether the exponential trend holds on real-world tasks. We include results from three supplementary external validity experiments.
First, we score our HCAST and RE-Bench tasks against a list of 16 “messiness” factors that aim to capture some systematic differences between our tasks and hard“real-world” tasks. Some examples include whether the task is resource limited, novel, or involves a dynamic environment (Section D.4). Controlling for task length, we find models perform worse on tasks that have higher messiness scores. Notably, trends in AI agent performance over time are similar for the lower and higher messiness subsets of our tasks (see Figure 9). In particular, we find no evidence of plateaus in performance trends specific to our higher messiness subset.
Second, we replicate our methods on SWE-bench Verified (Section 6.3), which includes human difficulty annotations. We find that the exponential trend still holds (Figure 11), albeit with an even shorter doubling time, possibly because SWE-bench Verified difficulty annotations, meant to represent high-context maintainer time, may differentially underestimate how long human contractors take to perform easier SWE-bench tasks.
Third, we measure AI agent performance on a small set of our internal pull requests (PRs) (Section 6.4). We find large differences in the speed at which different human groups complete the internal PR tasks, with contractors taking 5-18x longer to fix issues than repo maintainers. When using contractor (rather than maintainer) time-to-complete as a measure of task length, model time horizons derived from combined performance on SWAA, HCAST and RE-Bench are compatible with AI agent performance on internal PRs.
Our supplementary experiments find little evidence of performance trends being slower on the somewhat more realistic tasks we tested, but do not rule out the possibility that trends are meaningfully slower on the distribution of tasks required to automate software engineer jobs. These experiments were designed to detect large trend differences and constant factor shifts in model performance. We also find evidence that AI agent time horizons can differ by a large factor depending on the task domain and reference human population.
We conclude by discussing implications for AI capabilities forecasting (Section 8). Naively extrapolating the trend in horizon length implies that AI will reach a time horizon of >1 month (167 work hours) between late 2028 and early 2031 (Figure 12). However, extrapolation is affected by both external validity concerns and future changes in the growth rate. We discuss several possible factors that could either speed up or slow down the future trend in Section 7.2.
2 Related work
2.1 Agent and capability benchmarks
The evaluation of AI capabilities has evolved significantly from single-task benchmarks to complex, multi-step evaluations designed to assess agent-like behavior. While traditional benchmarks such as GLUE [9], SuperGLUE [10], and MMLU [11] have provided valuable insights into language model performance, they primarily measure static knowledge rather than the dynamic problem-solving capabilities essential for real-world applications. Recent work has developed more complex agent benchmarks. AgentBench [12] evaluates agents across diverse environments including web browsing, coding, and game playing. MLAgentBench [13] focuses specifically on machine learning research tasks, while ToolBench [14] assesses tool use capabilities. The recent ZeroBench [15] involve difficult reasoning, but in the context of visual puzzles rather than economically valuable tasks. Other noteworthy benchmarks include GAIA [16], which evaluates reasoning across multiple modalities, and BIG-bench [17], which contains hundreds of diverse tasks including many requiring multi-step reasoning.
Software engineering has emerged as a particularly informative domain for evaluating AI capabilities. HumanEval [18] and MBPP [19] provide programming challenges of varying complexity, while more complex benchmarks like SWE-bench [20] and APPS [21] test more sophisticated programming abilities. We use SWE-bench Verified [7]’s human time estimates for task completion in our work. RE-Bench [2], which we incorporate in our dataset, evaluates models on complex research engineering tasks that may require hours of human effort and compares AI performance to human machine learning engineers.
2.2 Forecasting AI progress
Several recent efforts have been made to contextualize AI benchmark performance. One such effort is the annual AI Index Report [6], which tracks performance across various benchmarks, including the date at which models achieved human-level performance. Murray et al. [27] had cybersecurity experts relate AI performance on Cybench [28] to the ability to autonomously develop malware. Phuong et al. [3] commissioned professional forecasters to predict whether AI ranks amongst top public concerns by 2030, conditioned on benchmark performance. These efforts generally lack a unified, quantitative metric for cross-benchmark comparison.
2.3 Psychometric methods and Item Response Theory
Our methodological approach draws inspiration from psychometric testing, particularly Item Response Theory (IRT) [32], which models the relationship between latent traits (such as ability) and observed responses to test items. In traditional IRT, item difficulty is a parameter in a logistic model predicting response correctness based on respondent ability. Our approach inverts this, using task completion time (a proxy for difficulty) to predict AI performance. Our methodology also relates to difficulty estimation techniques in educational testing [33], where multiple metrics including completion time are used to estimate the difficulty of tasks. IRT has been applied to machine learning classifiers by Martínez-Plumed et al. [34], and was used to design efficient benchmarks in Song and Flach [35].
@@ 3.1 Task suite / dataset
Our tasks are made up of three distinct task suites:
@@
    A subset of HCAST [8]: 97 diverse software tasks ranging from 1 minute to around 30 hours. [^5] Our results also include one task from GAIA [16], and five tasks involving writing code that is robust to an adversary, which are not included in HCAST.
@@
All tasks are automatically scored with a continuous score or binary threshold; details of how we normalize and process scores are given in Section 4.1. As with most benchmarks, our three task suites were also designed to isolate a specific unit of work that can be reliably accomplished within a time limit. This usually means that the tasks require much less context than the average task in the middle of a larger project. [^6] We define context as information that experienced employees use to complete a task which is not explicitly in the task description or possessed by most external experts. For example, when fixing a bug in a software package, the package’s maintainer may use their experience with past bugs in the same package to guess at the bug’s cause, fluency with the codebase to find the bug, and knowledge of their organization’s priorities to decide whether to apply a quick patch or a more thorough fix. We discuss possible effects of context in Section 6. We confirmed that all tasks were doable given the instructions provided by having humans successfully complete each of our tasks at least once. [^7] As many of these attempts were done by in-house staff or were done using different methodology, we exclude these attempts from our human baseline numbers.
3.1.1 HCAST tasks
Each task is automatically scored between 0 and 1, with higher scores indicating better performance. Many tasks only return scores of 0 or 1, but for tasks with continuous scoring, we manually define a success threshold score, which we use in some of our analysis to binarize agent scores. [^8] A small subset of these tasks is publicly available at https://github.com/METR/public-tasks. (We do not share the content of most tasks to reduce the likelihood of AI systems accidentally or intentionally being trained on them.)
3.1.2 RE-Bench suite
3.1.3 Software atomic actions (SWAA) suite
Example SWAA task (file selection) In contrast to other simple benchmarks like LAMBADA [36] or GSM8K [37], which test skills not directly applicable to software engineering, the SWAA set represents actions that are needed in both software engineering work and our longer tasks. SWAA consists of 5 task families, three representing common decisions, one for code completion, and one for math; see Appendix A.1 for more details. Development of SWAA tasks was blind to AI agent performance; that is, all tasks were written before seeing AI attempts, and elicitation (development of the few-shot prompt used for evaluation) was carried out on a separate development task suite.
3.1.4 Examples of tasks of varying lengths
Table 1 contains example tasks. Tasks that take less than one minute measure knowledge which would be needed for professional software engineering but do not require agency. At around one minute, tasks become similar to components of economically valuable work but are not entire projects (e.g. answering a basic question on Wikipedia). At about ten minutes, tasks become the difficulty of approximately the easiest meaningful step of a real software project (e.g. configuring a common open-source package). The shortest tasks which could reasonably be said to be standalone economically valuable projects take around one hour (e.g. writing a script to transform data from one format into another) and by eight hours tasks represent meaningfully valuable software projects (e.g. speeding up a python script by implementing custom CUDA kernels). Many more examples of these tasks are provided in Wijk et al. [2] and Rein et al. [8].
@@ 3.1.4 Examples of tasks of varying lengths
Table 1: Example tasks of differing durations. More examples can be found in Rein et al. [8] and Wijk et al. [2].
@@ 3.2 Baselining
In order to ground AI agent performance, we also measure the performance of multiple human “baseliners” on most of our tasks and recorded the duration of their attempts. In total, we use over 800 baselines totaling 2,529 hours, of which 558 baselines (286 successful) come from HCAST and RE-Bench, and 249 (236 successful) from the shorter SWAA tasks.
Baseliner skill and experience
@@
Our baseliners are skilled professionals in software engineering, machine learning, and cybersecurity, with the majority having attended world top-100 universities. They have an average of about 5 years of relevant experience, with software engineering baseliners having more experience than ML or cybersecurity baseliners.
3.2.1 HCAST tasks
We use existing baselines collected as part of HCAST. These baselines are collected from domain professionals with relevant experience in software engineering, ML, and cybersecurity. Baseliners work in the same environment as agents, using Vivaria, [^9] Our open source platform for language model agent evals: https://vivaria.metr.org/. with their screens and audio recorded for manual review, to prevent cheating. They are incentivized with bonuses for successful completion and for completing tasks faster than other baseliners. After screening out failed attempts and those with issues (such as using disallowed AI tools), we use 286 successful baselines from approximately 460 total attempts. Task durations are calculated using the geometric mean of successful baselines, with manual estimates for tasks lacking successful baselines. [^10] Non-baseline estimates were based on information including the length of QA runs that did not follow strict baseline conditions, and the length of similar tasks with successful baselines.
3.2.2 RE-Bench
For RE-Bench, we used the baselines from the RE-Bench paper. As baseliners were instructed to achieve the best performance for each task, we consider the task duration of each of these 6 tasks as 8 hours, and instead use the mean score achieved by baseliners who spent between 7 and 9 hours to convert raw score into success threshold.
3.2.3 SWAA
Unlike HCAST and RE-Bench, which were baselined by external contractors, SWAA is baselined by METR employees with relevant expertise using a custom webapp that enables more accurate timing. Because these tasks are intended to be a single step and exclude context acquisition, the timer for SWAA tasks ends as soon as the user chooses a response. For decision based tasks, only one selection is allowed to avoid random guessing; for fill in the blank tasks, baseliners can try multiple times until getting the answer correctly or opting to skip. We baselined each decision-based task 4 times and each fill-in-the-blank style task 3 times.
Suite Time Estimate Source Number of Tasks
HCAST Estimate 21
Baseline 76
RE-Bench Estimate 0
Baseline 7
SWAA Estimate 0
Baseline 66
Table 2: The source of our time estimates by task suite. In total, 148 of our 169 tasks have human baselines, but we rely on researcher estimates for 21 tasks in HCAST.
3.3 Evaluating AI agent performance on task suites
Most models we evaluated in this paper were models we had previously evaluated and therefore could reuse our scaffolding. We also included the earlier frontier models gpt-3.5-turbo-instruct, davinci-002 (GPT-3), and GPT-2. Full information on the models and agents used can be found in Appendix B.3.
3.3.1 Agent scaffolds
Most AI models were evaluated with modular-public—our basic agent scaffold. [^11] Code for this agent scaffold can be found at https://github.com/poking-agents/modular-public. This scaffold provides the model with Python and Bash commands and some very simple context management to keep the input within the context window length of the LM. We used a slightly different scaffold for o1-preview and o1, because they seemed to struggle with tool use, responding to environmental feedback, and generally acting as an agent. These are described further in Appendix B.3.
Figure 3: Average task success rate across our entire combined suite, for each model. As with all of the results reported in the main body of this work, to reduce the influence of large task families, we weight each task by the inverse square root of the number of tasks in the family it belongs to.
3.3.2 Results
@@
We performed 8 runs [^12] This number is approximate, because a small number of runs failed due to internal infrastructure issues. per agent/task pair and report the average results in Figure 3. As with most benchmarks, we notice a strong upwards trend over time, with recent models completing approximately 50% of all tasks, while earlier models perform substantially worse. We note that there is substantial correlation between the tasks that models can complete (Figure 22), with average correlation of approximately 0.73.
3.3.3 Model success rate vs baseline time
@@
There is a negative correlation between the time it takes a human baseliner to complete a task and the average success rate (across all models) on the task. This decrease in success rate over length (Figure 4) is well-fit by an exponential model ($R^{2}\approx 0.83$ when regressing model success rate against the logarithm of human time-to-complete). Notably, the correlation of model success with log human time (0.91) is higher than the average correlation between models (0.73, see Figure 22).
We sanity check this fit by examining the human-time-to-complete of tasks that earlier versus later models can complete. As expected, pre-2023 models like GPT-2 or GPT-3 can complete tasks requiring only writing a few words, but fail all tasks above 1 minute. In contrast, recent frontier models such as Claude 3.5 Sonnet (new) and o1 can complete some tasks that take human baseliners more than 4 hours (Figure 5).
4 Computing time horizon
To calculate a more intuitive metric for AI capabilities progress, we convert the performance of each model on our tasks to an estimate of their task completion time horizons.
4.1 From raw data to time horizon
First, the agent performance on each task is converted to a binary value (success or failure). Many tasks are naturally binary, including all SWAA tasks and the majority of HCAST tasks. Some tasks are continuously scored; these are binarized via a task-specific threshold. For example, if the task is to minimize the loss of a model, runs are binarized based on whether they achieved a loss below some threshold. Note that binarization particularly affects more challenging tasks such as RE-Bench, since current frontier models’ partial progress is usually binarized to 0 (as they are below human level on these tasks).
Figure 4: Model success rates are negatively correlated with how much time it takes a human to complete the task. ($y=-0.07x+0.66$, $R^{2}:0.83$)
Figure 5: Success rates of all models on our test suite, showing the computation of time horizon as predicted 50% success rate time. The logistic fit is fairly good, though there is a jump in success rate between <1 minute tasks and >1 minute tasks, which corresponds to the boundary between SWAA and HCAST tasks.

### v3

# Measuring AI Ability to Complete Long Software Tasks

@@ Abstract
Despite rapid progress on AI benchmarks, the real-world meaning of benchmark performance remains unclear. To quantify the capabilities of AI systems in terms of human capabilities, we propose a new metric: 50%-task-completion time horizon, the time humans typically take to complete tasks that AI models can complete with 50% success rate. We first timed humans with relevant domain expertise on a combination of RE-Bench, HCAST, and 66 novel shorter tasks. On these tasks, current frontier AI models such as o3 have a 50% time horizon of around 110 minutes. Furthermore, frontier AI time horizon has doubled approximately every seven months since 2019, though the trend may have accelerated since 2024. The increase in AI models’ time horizons seems to be primarily driven by greater reliability, ability to adapt to mistakes, logical reasoning, and capacity for tool use. We discuss the limitations of our results—including their degree of external validity—and the implications of increased autonomy for dangerous capabilities. If these results generalize to real-world software tasks, extrapolation of this trend predicts that within 5 years, AI systems will be capable of automating many software tasks that currently take humans a month.
@@ 1 Introduction
In the last five years, frontier AI systems have undergone a dramatic transformation in capabilities, evolving from basic text generation [37] to autonomously executing complex multi-hour machine learning research projects [46]. Sufficiently capable AIs could perform dangerous, highly complex actions like autonomous development of chemical, biological, radiological or nuclear weapons (CBRN) and self-replication and adaptation outside human control [34]. Understanding AI capabilities helps inform the development of safety guardrails as systems become increasingly powerful. In particular, many frontier AI developers have committed to using measures of specific AI capabilities to determine the necessary risk mitigations for their frontier AI systems. Robust benchmarks that can accurately track and forecast AI capabilities thus form the foundation for responsible AI governance and risk mitigation.
@@
However, existing benchmarks face several key limitations. First, they often consist of artificial rather than economically valuable tasks. Second, benchmarks are often adversarially selected for tasks that current models struggle with compared to humans, [^1] For example, HellaSwag [50] and Humanity’s Last Exam [33] were both generated by adversarially filtering problems against the best performing language models available at the time. biasing the comparison to human performance. Most critically, individual benchmarks saturate increasingly quickly [24], and we lack a more general, intuitive, and quantitative way to compare between different benchmarks, [^2] SWE-bench Verified [7] does come with human-estimated task completion times. We use SWE-bench Verified tasks and accompanying time estimates to validate our main result in Section 4.1. which prevents meaningful comparison between models of vastly different capabilities (e.g., GPT-2 versus Claude 3.7 ). As a consequence, while the last few years have seen dramatic increases in AI performance on many individual benchmarks, understanding the progress of AI capabilities in general has required estimating the qualitative difficulty of the latest benchmarks AI systems can pass.
We propose tracking AI progress over time using the task completion time horizon: the duration of tasks that models can complete at a certain success probability, providing an intuitive measure of real-world capability compared to humans. As models may not reliably complete all tasks of a given length, we operationalize this by measuring the X%-(task completion) time horizon–the length of tasks that models can complete approximately X% of the time.
We prototype this methodology using three datasets designed to capture skills required for research or software engineering (Section 2.1), totaling 170 tasks with a wide range of difficulty: HCAST [38], RE-Bench [46], and Software Atomic Actions (SWAA), a new suite of shorter software tasks that can measure pre-2023 models (Appendix B.1.3). Using skilled human baseliners, we estimate the duration that a domain knowledgeable human (without task-specific context) takes to complete these tasks (Section 2.2). We evaluate the performance of 12 frontier models from 2019 to 2025 on these tasks (Section 2.3). Using methodology inspired by human psychometric studies, we then estimate the duration of tasks that models can complete with 50% success rate—the 50% time horizon (Section 3.1).
We find that the 50% time horizon has been growing exponentially from 2019–2025 on these tasks, with a doubling time of approximately seven months (Figure 1). We compare our main result with our exploratory results on non-SWAA tasks, finding that the 2023–2025 horizon growth rate is about 20% faster than the 2019–2025 rate (Section F.1). We also measure the 80% time horizon of models (Figure 17) and find a similar trend, though horizons are roughly 5x shorter. This progress appears to be driven by several key factors: improved logical reasoning, better tool use, and greater reliability and self-awareness in task execution (Section 3.3). We also note several limitations of current systems—notably, performance is much lower on less structured, “messier” tasks (Section F.2).
Since our tasks do not perfectly represent the average segment of intellectual labor by researchers and software engineers, this raises the question of external validity (Section 4): whether the exponential trend holds on real-world tasks. We include three supplementary external validity experiments, which find little evidence of performance trends being slower on the somewhat more realistic tasks we tested, but do not rule out the possibility that trends are meaningfully slower on the distribution of tasks required to automate software engineer jobs. We also find evidence that AI agent time horizons can differ by a large factor depending on the task domain and reference human population.
We conclude by discussing implications for AI capabilities forecasting (Section 5). Naively extrapolating the trend in horizon length implies that AI will reach a time horizon of >1 month (167 work hours) between mid-2028 and mid-2031 (Figure 6). However, extrapolation is affected by both external validity concerns and future changes in the trend.
@@
1.1 Related Work
Here, we briefly discuss related work. See Appendix A for an expanded discussion.
Agentic AI capability benchmarks
Many recent benchmarks aim to measure the capability of AI systems to act as agents in general, including AgentBench [22], ToolBench [13], GAIA [26], TheAgentCompany [47], the Berkeley Function Calling Leaderboard [48]. In this work, we study AI agent performance on RE-Bench[46], HCAST [38], and SWE-Bench Verified [7], which focus on machine learning and software engineering tasks. Other similar benchmarks include MLAgentBench [17], MLEBench [5], and DSBench [20]. In concurrent work, Miserendino et al. [27] create the SWE-Lancer benchmark, consisting of 1400 freelance software engineering tasks gathered from Upwork with real-world payouts totalling above \$1 million.
AI capability forecasting
This work builds on previous efforts to forecast the capabilities of AI systems over time, including efforts to contextualize benchmark performance using human experts Phuong et al. [34]; Murray et al. [28] and efforts to forecast performance on particular benchmarks Owen [31]; Pimpale et al. [35]. Contextualizing AI capabilities using human time horizon was proposed by Ngo [29], and similar concepts were discussed in Carlsmith [4] and Cotra [9].
Human psychometric methods
Our methodology draws inspiration from human psychometric testing, especially Item Response Theory [3; 10; 23].
@@ 2.1 Task suite / dataset
Our tasks comprise three distinct task suites:
@@
    A subset of HCAST [38]: 97 diverse software tasks ranging from 1 minute to 30 hours. [^3] Our results also include one task from GAIA [26], and five tasks involving writing code that is robust to an adversary.
@@
All tasks are automatically scored with a continuous score or binary threshold; details of how we normalize and process scores are given in Section 3.1, and other task suite details in Appendix B.
Example tasks
See Table 1 for five example tasks. Tasks under 1 minute measure knowledge relevant to professional software engineering. At 10 minutes, tasks can include the easiest meaningful step of a real software project (e.g. configuring a common open-source package). The shortest tasks that could be standalone economically relevant projects take around one hour; by eight hours tasks represent valuable software projects. Many more examples of these tasks are available in Wijk et al. [46] and Rein et al. [38].
@@
Table 1: Example tasks for different durations; see Rein et al. [38] and Wijk et al. [46] for more examples.
@@ 2.2 Baselining
In order to ground AI agent performance, we also measure the performance of multiple human “baseliners” on most tasks and recorded the duration of their attempts. In total, we use over 800 baselines totaling 2,529 hours, of which 558 baselines (286 successful) come from HCAST and RE-Bench, and 249 (236 successful) from the shorter SWAA tasks. 148 of the 169 tasks have human baselines, but we rely on researcher estimates for 21 tasks in HCAST.
@@
Our baseliners are skilled professionals in software engineering, machine learning, and cybersecurity, with the majority having attended world top-100 universities. They have an average of about 5 years of relevant experience, with software engineering baseliners having more experience than ML or cybersecurity baseliners. For more details about baselines, see Appendix C.1.
2.3 Results
@@
We measure 12 frontier (and 4 near-frontier) models released between 2019 and 2025; full information on the models and scaffolds used can be found in Appendix C.3. We perform 8 runs [^4] This number is approximate, because a small number of runs failed due to internal infrastructure issues. per agent/task pair and report the average results in Figure 3. There is a strong upwards trend over time, with recent models completing approximately 50% of all tasks, while earlier models perform substantially worse. We find substantial correlation between the tasks that models can complete (Figure 22), with average correlation of approximately 0.73.
Figure 3: Left: Average task success rate across the entire combined suite, for each model. As with all of the results reported in the main body of this work, to reduce the influence of large task families, we weight each task by the inverse square root of the number of tasks in the family it belongs to. Right: Model success rates are negatively correlated with how much time it takes a human to complete the task. (y=0.07x+0.66y=-0.07x+0.66, $R^{2}:0.83$)
@@
There is a negative correlation between the time it takes a human baseliner to complete a task and the average success rate (across all models) on the task. This decrease in success rate over length (Figure 3) is well-fit by an exponential model ($R^{2}\approx 0.80$ when regressing model success rate against the logarithm of human time-to-complete). As expected, more recent models can complete longer tasks, with Claude 3.7 Sonnet and o3 completing some tasks that take human baseliners >4 hours (Figure 4).
3 Time horizon
3.1 Computing time horizon